# Invariant Coordinate Selection and Fisher discriminant subspace beyond the case of two groups

### v1

@@ Appendix C.2.   The case of $k$ groups
Subsection 4.2 generalizes the context to a number $k$ of groups and focuses on the notion of “thresholds”. Figure 6 is an illustration of the thresholds found using three different setups, for 2 to 10 groups. The threshold values are described in Table 5.
Table 5: Thresholds computed for each setup
k Setup 1 Setup 2 Setup 3
3 0.18 0.2 0.171
4 0.155 0.161 0.151
5 0.135 0.138 0.133
6 0.12 0.121 0.119
7 0.107 0.108 0.107
8 0.097 0.097 0.097
9 0.089 0.089 0.088
10 0.082 0.082 0.081
@@
Table 6: Grids used in Setups 1, 2, and 3 for $k=4$
Setup 1 Setup 2 Setup 3
$\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ $\alpha_{4}$ $\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ $\alpha_{4}$ $\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ $\alpha_{4}$
0.25 0.25 0.25 0.25 0.25 0.175 0.25 0.325 0.25 0.05 0.25 0.45
0.249 0.25 0.25 0.251 0.249 0.175 0.25 0.326 0.249 0.05 0.25 0.451
0.248 0.25 0.25 0.252 0.248 0.175 0.25 0.327 0.248 0.05 0.25 0.452
0.001 0.25 0.25 0.499 0.001 0.175 0.25 0.574 0.001 0.05 0.25 0.699
@@
The second setup (Setup 2 in Figure 6) uses a grid based on the results obtained with the first setup. From the Setup 1 grid, the value of the second group proportion is updated. It is set to the threshold found in Setup 1 to which 2% is added. The proportion of the first group is the same, ranging from 0.001 to $1/k$ with a step of 0.001. The intermediate groups have a proportion of $1/k$ only for group 3 through $k-1$, since the proportion of the second group is already set. The value of the last group proportion is also updated, but it is still equal to one minus the other group weights. This procedure ensures that the sum of each row is exactly one. An example of the grid obtained with the second setup when $k=4$ is shown in Table 6. However, it requires at least three groups: one that is set to the first threshold plus 2%, and two groups with variable proportions. For each configuration of group weights, the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ are calculated. Initially, the first group is not necessarily the smallest one; it could be the second group. However, the value of the second group proportion is set such that it is above the threshold, implying that all eigenvalues are initially below one. Thus, the threshold of Setup 2 is, as in Setup 1, the first value of the first group weight for which one eigenvalue exceeds one. The concept in this procedure revolves around examining the impact of a group that is positioned near the threshold to determine if this proximity affects it. This process is repeated for a number of groups ranging from 3 to 10.
@@
Finally, the third procedure (Setup 3 in Figure 6) uses a grid similar to Setup 2. The second group proportion is no longer dependent on the threshold of Setup 1 but is set to 0.05. Again, the last group proportion is updated to one minus the combined weights of the other groups. All other group proportions remain the same as in Setups 1 and 2. Table 6 provides an example of the third setup grid with $k=4$. As for the grid in Setup 2, it requires at least three groups: one for the fixed proportion and two to vary. The objective of this procedure is to include a small group initially, specifically below the threshold for $k$ from 3 to 10. This implies that, from the initial point, there is already an eigenvalue greater than one. Thus, the threshold of Setup 3 is the first value of the first group weight for which two eigenvalues exceed one. This is aimed at identifying the point where the second eigenvalue surpasses one. This setup is repeated for 3 to 10 groups.
Appendix D.   Proof of Proposition 2
Proof.
@@ Appendix D.   Proof of Proposition 2
By adapting the calculations in Appendix A (with the same notations) to the Gaussian mixture model (9), we obtain
$\mathrm{COV}=\begin{bmatrix}\beta_{11}&0\\ 0&I_{p-1}\end{bmatrix},\quad\mathrm{COV}^{-1}=\begin{bmatrix}b_{11}&0\\ 0&I_{p-1}\end{bmatrix},\quad\text{and}\quad\mathrm{COV}_{4}$ $=\begin{bmatrix}a_{11}&0\\ 0&I_{p-q},\end{bmatrix}$
@@
where $\beta_{11}=1+\sum_{j=1}^{3}\alpha_{j}(t_{j1}^{\mbox{c}})^{2}$, $b_{11}=\beta_{11}^{-1}$ and $a_{11}=\frac{1}{p+2}\Bigg{[}b_{11}\mathbb{E}[(x^{\mbox{c}}_{1})^{4}]+(p-1)\mathbb{E}[(x^{\mbox{c}}_{1})^{2}]\Bigg{]}$. We have
$\mathrm{COV}^{-1}\mathrm{COV}_{4}=\begin{bmatrix}b_{11}a_{11}&0\\ 0&I_{p-1}.\end{bmatrix}$
@@
For the mixture (6), we know from Theorem 4 in Tyler et al. (2009) and using the computations of Subsection 2.4, that at least $p-1$ eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ are equal to one. As for the remaining eigenvalue, using the computations of Subsection 2.4 and Mathematica Wolfram Research, Inc. (2022), we can prove that it is equal to one if and only if
$p_{\alpha_{1},\alpha_{2}}(t_{11},t_{12})=0,$ (10)
@@
$p_{\alpha_{1},\alpha_{2}}(t_{11},t_{12})=\alpha_{1}(-1+7\alpha_{1}-12\alpha_{1}^{2}+6\alpha_{1}^{3})t_{11}^{4}+4\alpha_{1}\alpha_{2}(1-6\alpha_{1}+6\alpha_{1}^{2})t_{11}^{3}t_{12}+6\alpha_{1}\alpha_{2}(1-2\alpha_{2}+$
$\alpha_{1}(-2+6\alpha_{2}))t_{11}^{2}t_{12}^{2}+4\alpha_{1}\alpha_{2}(1-6\alpha_{2}+6\alpha_{2}^{2})t_{11}t_{12}^{3}+\alpha_{2}(-1+7\alpha_{2}-12\alpha_{2}^{2}+6\alpha_{2}^{3})t_{12}^{4}.$
Since $t_{21}\neq 0$, we have

### v2

@@ Appendix B.2.   The case of $k$ groups
Subsection 4.2 generalizes the context to a number $k$ of groups and focuses on the notion of “thresholds”. Fig. 6 is an illustration of the thresholds found using three different setups, for 2 to 10 groups. The threshold values are described in Table 1.
Table 1: Group proportion thresholds computed for different values of the number of groups $k$ and three setups. Setup 1: all group proportions initially equal. Setup 2: $\alpha_{2}$ near the Setup 1 threshold. Setup 3: $\alpha_{2}$ equal to 5%.
  k Setup 1 Setup 2 Setup 3 3 0.18 0.2 0.171 4 0.155 0.161 0.151 5 0.135 0.138 0.133 6 0.12 0.121 0.119 7 0.107 0.108 0.107 8 0.097 0.097 0.097 9 0.089 0.089 0.088 10 0.082 0.082 0.081
@@
Table 2: Group proportions ($\alpha_{1},\ldots,\alpha_{4}$) used in Setups 1, 2, and 3 for $k=4$ groups. Setup 1: all group proportions initially equal. Setup 2: $\alpha_{2}$ near the Setup 1 threshold. Setup 3: $\alpha_{2}$ equal to 5%.
  Setup 1 Setup 2 Setup 3 $\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ $\alpha_{4}$ $\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ $\alpha_{4}$ $\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ $\alpha_{4}$ 0.25 0.25 0.25 0.25 0.25 0.175 0.25 0.325 0.25 0.05 0.25 0.45 0.249 0.25 0.25 0.251 0.249 0.175 0.25 0.326 0.249 0.05 0.25 0.451 0.248 0.25 0.25 0.252 0.248 0.175 0.25 0.327 0.248 0.05 0.25 0.452 0.001 0.25 0.25 0.499 0.001 0.175 0.25 0.574 0.001 0.05 0.25 0.699
@@
The second setup (Setup 2 in Fig. 6) uses a grid based on the results obtained with the first setup. From the Setup 1 grid, the value of the second group proportion is updated. It is set to the threshold found in Setup 1 to which 2% is added. The proportion of the first group is the same, ranging from 0.001 to $1/k$ with a step of 0.001. The intermediate groups have a proportion of $1/k$ only for group 3 through $k-1$, since the proportion of the second group is already set. The value of the last group proportion is also updated, but it is still equal to one minus the other group weights. This procedure ensures that the sum of each row is exactly one. An example of the grid obtained with the second setup when $k=4$ is shown in Table 2. However, it requires at least three groups: one that is set to the first threshold plus 2%, and two groups with variable proportions. For each configuration of group weights, the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ are calculated. Initially, the first group is not necessarily the smallest one; it could be the second group. However, the value of the second group proportion is set such that it is above the threshold, implying that all eigenvalues are initially below one. Thus, the threshold of Setup 2 is, as in Setup 1, the first value of the first group weight for which one eigenvalue exceeds one. The concept in this procedure revolves around examining the impact of a group that is positioned near the threshold to determine if this proximity affects it. This process is repeated for a number of groups ranging from 3 to 10.
@@
Finally, the third procedure (Setup 3 in Fig. 6) uses a grid similar to Setup 2. The second group proportion is no longer dependent on the threshold of Setup 1 but is set to 0.05. Again, the last group proportion is updated to one minus the combined weights of the other groups. All other group proportions remain the same as in Setups 1 and 2. Table 2 provides an example of the third setup grid with $k=4$. As for the grid in Setup 2, it requires at least three groups: one for the fixed proportion and two to vary. The objective of this procedure is to include a small group initially, specifically below the threshold for $k$ from 3 to 10. This implies that, from the initial point, there is already an eigenvalue greater than one. Thus, the threshold of Setup 3 is the first value of the first group weight for which two eigenvalues exceed one. This is aimed at identifying the point where the second eigenvalue surpasses one. This setup is repeated for 3 to 10 groups.
Appendix C.   Proof of Proposition 2
Proof:.
@@ Appendix C.   Proof of Proposition 2
By adapting the calculations in Appendix A (with the same notations) to the Gaussian mixture model (6), we obtain:
$\mathrm{COV}=\begin{bmatrix}\beta_{11}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-1}\end{bmatrix},\quad\mathrm{COV}^{-1}=\begin{bmatrix}b_{11}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-1}\end{bmatrix},\quad\mathrm{COV}_{4}$ $=\begin{bmatrix}a_{11}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-1},\end{bmatrix},$
@@
where $\beta_{11}=1+\sum_{j=1}^{3}\alpha_{j}(t_{j1}^{\mbox{c}})^{2}$, $b_{11}=\beta_{11}^{-1}$ and $a_{11}=\frac{1}{p+2}\Bigg{[}b_{11}\mathbb{E}[(x^{\mbox{c}}_{1})^{4}]+(p-1)\mathbb{E}[(x^{\mbox{c}}_{1})^{2}]\Bigg{]}$. We have:
$\mathrm{COV}^{-1}\mathrm{COV}_{4}=\begin{bmatrix}b_{11}a_{11}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-1}\end{bmatrix}.$
@@
For the mixture (6), we know from Theorem 4 in Tyler et al. (2009) and using the computations of Subsection 2.4, that at least $p-1$ eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ are equal to one. As for the remaining eigenvalue, using the computations of Subsection 2.4 and Mathematica Wolfram Research, Inc. (Champaign, IL, 2022), we can prove that it is equal to one if and only if
$p_{\alpha_{1},\alpha_{2}}(t_{11},t_{12})=0,$
@@
$p_{\alpha_{1},\alpha_{2}}(t_{11},t_{12})$ $=\alpha_{1}(-1+7\alpha_{1}-12\alpha_{1}^{2}+6\alpha_{1}^{3})t_{11}^{4}+4\alpha_{1}\alpha_{2}(1-6\alpha_{1}+6\alpha_{1}^{2})t_{11}^{3}t_{12}+6\alpha_{1}\alpha_{2}(1-2\alpha_{2}$
$+\alpha_{1}(-2+6\alpha_{2}))t_{11}^{2}t_{12}^{2}+4\alpha_{1}\alpha_{2}(1-6\alpha_{2}+6\alpha_{2}^{2})t_{11}t_{12}^{3}+\alpha_{2}(-1+7\alpha_{2}-12\alpha_{2}^{2}+6\alpha_{2}^{3})t_{12}^{4}.$
Since $t_{21}\neq 0$, we have:
@@
Appendix D.   Supplementary data
Supplementary material related to this article can be found online at (to be filled by the journal).